a) current through fuse at point "A"= _____________?

b) if the wires at B were connected, what then would the current through the fuse at point A equal?

Show all calculations

a) current through fuse at point
Answer #1

Given that,

V = 120 Volts;

We notive that all the resistors are connected in series, so the effective resistance would be simply the sum of all the resistances between the points A and B(there are 6 with 100 Ohm). So,

Req = 600 Ohm

From Ohm's law:

V = IR =. I = V/R

I = 120/600 = 0.2 A

The current in the wire is 0.2 A. And All the devices are connected in series and its a fact that the same current flows thorugh all the resistors connected in series.

Hence, current through the fuse point is I = 0.2 A.